# Balancing Act: Prioritization Strategies for LLM-Designed Restless Bandit Rewards

### v2

@@ Abstract
LLMs are increasingly used to design reward functions based on human preferences in Reinforcement Learning (RL). We focus on LLM-designed rewards for Restless Multi-Armed Bandits, a framework for allocating limited resources among agents. In applications such as public health, this approach empowers grassroots health workers to tailor automated allocation decisions to community needs. In the presence of multiple agents, altering the reward function based on human preferences can impact subpopulations very differently, leading to complex tradeoffs and a multi-objective resource allocation problem. We are the first to present a principled method termed Social Choice Language Model for dealing with these tradeoffs for LLM-designed rewards for multiagent planners in general and restless bandits in particular. The novel part of our model is a transparent and configurable selection component, called an adjudicator, external to the LLM that controls complex tradeoffs via a user-selected social welfare function. Our experiments demonstrate that our model reliably selects more effective, aligned, and balanced reward functions compared to purely LLM-based approaches.
1 Introduction
@@
Reward functions play a fundamental role in the generation of optimal policies for sequential decision-making via reinforcement learning. Previous work has shown that LLMs are an effective tool for designing reward functions that can be guided and customized via human language prompts (Ma et al. 2023; Cao et al. 2024; Kwon et al. 2023; Xie et al. 2024a; Ma et al. 2024; Yu et al. 2023; Hazra et al. 2024). We study the problem of designing high-quality reward functions aligned with human preference prompts in the context of multiagent optimization and planning problems. We present a transparent framework around LLMs that constructs effective, aligned, and balanced reward functions for complex human prompts.
@@
We study the reward design problem for restless multi-armed bandits (RMABs), a popular model in multiagent systems for sequentially allocating a limited number of resources to a set of agents (Whittle 1988; Niño-Mora 2023). In RMABs, each agent is represented by an individual Markov Decision Process including a reward function. By choosing these reward functions, one can control which agents are more or less likely to receive a resource. RMABs have been used in various domains such as machine maintenance (Abbou and Makis 2019), anti-poaching (Qian et al. 2016a), and healthcare (Ayer et al. 2019; Verma et al. 2023). In many of them, system organizers have evolving allocation priorities based on agents’ features that need to be incorporated into the resource allocation process (Deardorff et al. 2018; Verma et al. 2024). For instance, in a healthcare program, a healthcare worker might want to change the allocation policy to prioritize low-income beneficiaries who are at higher risk or older beneficiaries who have transportation barriers for healthcare access (Nelson et al. 2016; Syed, Gerber, and Sharp 2013) via the preference prompt: Prioritize low-income beneficiaries and older beneficiaries.
@@
Translating such human language prompts to effective and aligned reward functions is a general, non-trivial challenge in RL (Hazra et al. 2024; Cao et al. 2024; Ma et al. 2023), e.g., as it is unclear how changes in the reward influence the used policy downstream. However, the multiagent nature of the RMAB problem adds a new twist to the problem of reward design in RL: It becomes fundamentally multi-objective. Consider the above example prompt asking for the prioritization of two subpopulations. As these subpopulations may contain different agents, selecting a reward function will most likely involve trading off the interests of the low-income vs. older beneficiaries, making this a multi-objective problem. If this multi-objective nature is ignored, a selected reward function might heavily favor one of the two groups (e.g., leading to the allocation of many resources to low-income beneficiaries, and no resources to older ones). Another tradeoff that emerges because of the multiagent nature of the problem is between aligning with the preference prompt and impacting agents not referenced in the prompt: Changing the reward functions to align with a prompt can lead to drastic changes in the number of resources distributed to subpopulations not referenced in the user prompt (e.g., in the example prompt, the selected reward might lead to taking away all resources from well-educated beneficiaries). If not adequately balanced, this can cause strong side effects unintended by the user.
@@
We present a Social Choice Language Model (SCLM) that designs reward functions (as Python code) aligned with complex, multi-objective human language preferences; see Figure 2 for an overview of SCLM. Our pipeline separates the generation of candidate reward functions in the generator from the selection of one function in the adjudicator. For the generator, we use LLM-powered evolutionary search to generate a pool of reward functions (Behari et al. 2024). In the transparent and customizable adjudicator, we take a new social choice perspective to address the multi-objective nature of our problem: We create a scorer component that evaluates the quality of generated reward functions according to the different objectives (e.g., different prioritization requests). Subsequently, a social welfare function aggregates these “alignment scores” to select the best reward function. The user can select the social welfare function and thereby has additional control over the preferred trade-off between objectives, e.g., maximizing the summed vs. minimum alignment of all objectives. We demonstrate in our experiments that our SCLM leads to the selection of reward functions significantly better aligned with complex, multi-objective prompts. Moreover, we also show how it can be used to effectively mitigate the risks of using rewards designed from human prompts: unintended effects for other agents and the ineffective allocation of resources. SCLM combines the generative power of LLMs to design reward functions with the capabilities of social choice to handle multi-objective decision-making scenarios.
2 Related Works
@@
LLMs have emerged as a powerful tool to enhance RL. Recent work has used LLMs to generate reward functions based on natural language descriptions (Ma et al. 2023; Xie et al. 2024b; Yu et al. 2023). For instance, Goyal, Niekum, and Mooney (2019); Carta et al. (2022); Mirchandani, Karamcheti, and Sadigh (2021); Hazra et al. (2024) shape rewards by training an RL agent to learn and complete intermediate tasks guided by language, yet focusing on very different (non-multiagent) environments.
The work of Behari et al. (2024) is the first to present a Decision-Language Model for generating reward functions for RMABs from human prompts. The model performs a form of evolutionary search to find reward functions aligned with the given prompt in two interleaving phases: generation and reflection. In the generation phase, an LLM generates a set of reward functions. Based on reward function’s performances, in the reflection phase (Ma et al. 2023; Shinn et al. 2023), the LLM selects the function best aligned with the prompt. This function is then included in the prompt for the next generation phase or returned. In contrast to our work, DLM mixes generation with selection and does not explicitly account for the multi-objective nature of the reward design problem. Furthermore, in contrast to our work, they focus on small RMAB instances ($\sim 20$ arms). Throughout the paper, we will use a slightly modified variant of DLM adjusted to our setting (see Section C.2) as a baseline.
LLMs can also be integrated into model-based and explainable RL, serving as world model simulators (Lin et al. 2023) or policy interpreters (Das, Chernova, and Kim 2023). They can also act as information processors, effectively managing environment data and task instructions (Basavatia, Murugesan, and Ratnakar 2024; Song et al. 2023) and as direct decision makers (Li et al. 2022) to improve the sample efficiency of offline RL. While this area of research is rapidly evolving, there remains a significant gap in addressing multi-objective, multi-agent scenarios within LLM-enhanced RL, which we begin to address in this paper.
@@
Research on MORL focuses on learning policies that maximize (and balance between) multiple objective functions, typically via scalarizing the objectives into a single reward function (Moffaert, Drugan, and Nowé 2013) or approximating the Pareto front (Roijers et al. 2013; Van Moffaert and Nowé 2014). In the context of multiagent systems, MORL has been used as a method to ensure the fair treatment of the individual agents (Jiang and Lu 2019; Zimmer et al. 2021; Fan et al. 2023). Closest to ours from this line of work is the paper by Fan et al. (2023), who use ideas from the resource allocation literature to combine multiple objectives into a singular non-linear objective function. However, in contrast to our paper, their focus lies on policy learning for such non-linear objective functions. They neither consider reward design, human preference prompts, LLMs, nor RMABs.
@@
3 Preliminaries
An instance of Restless Multi-Armed Bandits (RMAB) is defined by a set of $N$ arms, a time horizon $T$, and a budget $K$. We also refer to arms as agents. Each arm $i\in[N]$ is an independently evolving MDP with state space $\mathcal{S}_{i}$, actions $\mathcal{A}_{i}=\{0,1\}$, transition function $P_{i}:\mathcal{S}_{i}\times\mathcal{A}_{i}\times\mathcal{S}_{i}\to\mathbb{R}_{\geq 0}$, and reward function $R_{i}:\mathcal{S}_{i}\to\mathbb{R}$. We refer to $1$ as the active action corresponding to pulling the arm (i.e., allocating a resource) and 00 as the passive action corresponding to not pulling the arm. We focus on the popular case where each MDP consists of two states, i.e., $\mathcal{S}_{i}=\{0,1\}$ for all $i\in[N]$, yet our methodology applies to MDPs with arbitrary state spaces. We refer to 00 as the bad and $1$ as the good state. For each step in which an agent is in the good state, they derive a utility of $1$, while they derive a utility of 00 in the bad state. Accordingly, agents’ default reward function $R^{*}$ is $R^{*}(s)=s$. We assume that there is a set of categorical features. Each arm is associated with a value of each feature. A global reward function is a reward function defined over features, which induces a reward function for each arm by plugging in its feature values (see Example 4.1).
In each step within the time horizon $T$, the planner observes the state of all arms and decides to pull a subset of at most $K$ arm. As solving the RMAB problem optimally is computationally intractable (Papadimitriou and Tsitsiklis 1994), we make use of the very popular state-dependent Whittle index (Whittle 1988; Niño-Mora 2023), which tries to quantify for each state of each arm the reward gain achieved from applying the active action to the arm in this state. In the Whittle index policy $\Pi$, in each step, we compute the Whittle index for each arm (based on its current state) and pull the arms with the $K$ highest Whittle indices. We will use it as the solution strategy in the following.
For a global reward function $R$, we write $\Pi(R)$ to denote the Whittle index policy for $R$, i.e., the Whittle index policy for the instance where each agent uses the function $R$ after plugging in their feature values as their reward. We refer to $\Pi(R^{*})$ as the default policy. To assess the quality of a global reward function $R$, we often consider the utility feature distribution for some feature $X$. This distribution shows for each value of the feature, the expected utility generated by arms with this feature value under the policy $\Pi(R)$ (see Figure 4(a) in Appendix B for an example).
4 Problem Statement & Challenges
@@
Figure 1: Tradeoffs between prioritization clauses.
@@
Our goal is to create a model that handles the tradeoffs posed by composite “multi-objective” prompts in a principled, transparent, and customizable fashion and outputs a single effective and fairly aligned global reward function.
5 Social Choice Language Model (SCLM)
Figure 2: Overview of SCLM. In step 1, the preference prompt is passed to the generator, which performs an evolutionary search to create a pool $\mathcal{R}$ of candidate reward functions. In step 2, these functions are passed to the adjudicator where a scorer model (e.g., the simulator or LLM scorer) computes the alignment scores. In step 3, a user-defined social welfare function selects a reward function based on the alignment scores.
@@ 5 Social Choice Language Model (SCLM)
5.1 Generator
Given a prompt, our generator creates a set of candidate reward functions (as Python code) via a variant of evolutionary search (Behari et al. 2024): We proceed in multiple steps. First, inputting the problem description, feature descriptions and the preference prompt, we ask an LLM to generate code for a reward function. We repeat this query $n_{p}$ times to obtain a set $\mathcal{R}$ of $n_{p}$ candidate reward functions. Afterwards, for each function $R\in\mathcal{R}$ we compute the utility feature distributions of the policy $\Pi(R)$ induced by the reward function $R$ on the given RMAB instance (via repeatedly simulating the policy on the instance). Then, the prompt and the set of candidate reward functions together with the associated utility feature distributions are passed to an LLM, which is asked to select the reward function $R^{\prime}$ from $\mathcal{R}$ best aligned with the prompt (Ma et al. 2023; Shinn et al. 2023). Now, we repeat the whole process, this time including the selected policy $R^{\prime}$ as a seed in the reward function generation prompts. Once we have executed the process $n_{r}$ times, we add all generated $n_{p}\cdot n_{r}$ candidate reward functions $R$ to the pool $\mathcal{R}$ (see Section C.2 for details).
5.2 Adjudicator
The adjudicator selects a reward function from a given pool of candidate reward functions returned by the generator. To handle the complex tradeoffs arising within composite prompts and the resulting multi-objective optimization problem, the adjudicator follows a social choice approach. Social choice is a discipline at the intersection of economics, philosophy, and mathematics and concerned with aggregating the potentially contradicting preferences of a set of voters into a fair compromise alternative from a given candidate set (Arrow, Sen, and Suzumura 2010; Moulin 2004). It thus provides a theoretically grounded and well-studied methodology for balancing competing interests. In our problem, we can interpret the reward functions as the candidates and the preference clauses in the prompt as the voters with their preferences over the candidates reflecting the reward function’s alignment with the clause. This view gives rise to the following strategy: Given a prompt $P=\{p_{1},p_{2},\dots,p_{\ell}\}$, we evaluate each reward function $R\in\mathcal{R}$ from the perspective of each preference clause $p_{i}$ by computing an (alignment) score $s_{i}(R)$. $s_{i}(R)$ measures the alignment of $\Pi(R)$ with preference clause $p_{i}$, i.e., how much the voter representing $p_{i}$ “likes” the candidate $R$. We describe at the end of this section how these scores are computed.

### v3

@@ Abstract.
LLMs are increasingly used to design reward functions based on human preferences in multiagent Reinforcement Learning (RL). We focus on LLM-designed rewards for Restless Multi-Armed Bandits, a framework for allocating limited resources among agents. In applications such as public health, this approach empowers grassroots health workers to tailor automated allocation decisions to community needs. In the presence of multiple agents, altering the reward function based on human preferences can impact subpopulations very differently, leading to complex tradeoffs and a multi-objective resource allocation problem. We are the first to present a principled method termed Social Choice Language Model for dealing with these tradeoffs for LLM-designed rewards for multiagent planners in general and restless bandits in particular. The novel part of our model is a transparent and configurable selection component, called an adjudicator, external to the LLM that controls complex tradeoffs via a user-selected social welfare function. Our experiments demonstrate that our model reliably selects more effective, aligned, and balanced reward functions compared to purely LLM-based approaches.
Key words and phrases:
Restless Bandits, Multi-Objective RL, Mobile Health
11footnotetext: Joint first author
1. Introduction
@@
Reward functions play a fundamental role in the generation of optimal policies for sequential decision-making via reinforcement learning. Previous work has shown that LLMs are an effective tool for designing reward functions that can be guided and customized via human language prompts Ma et al. (2023); Cao et al. (2024); Kwon et al. (2023); Xie et al. (2024a); Ma et al. (2024); Yu et al. (2023); Hazra et al. (2024). Focusing on optimization and planning scenarios, we study the problem of designing high-quality reward functions aligned with human preference prompts in a multiagent context, rendering the underlying problem inherently multi-objective. We present a transparent framework around LLMs that constructs effective, aligned, and balanced reward functions for complex human prompts.
@@
We study the reward design problem for restless multi-armed bandits (RMABs), a popular model in multiagent systems for sequentially allocating a limited number of resources to a set of agents Whittle (1988); Niño-Mora (2023). In RMABs, there are multiple, independently evolving agents, with each agent being represented by an individual Markov Decision Process including a reward function. By choosing these reward functions, one can control which agents are more or less likely to receive a resource. RMABs have been applied to multiagent problems in various domains such as machine maintenance Abbou and Makis (2019), anti-poaching Qian et al. (2016a), and healthcare Ayer et al. (2019); Verma et al. (2023). In many of them, system organizers have evolving allocation priorities based on agents’ features that need to be incorporated into the resource allocation process Deardorff et al. (2018); Verma et al. (2024). For instance, in a healthcare program, a healthcare worker might want to change the allocation policy to prioritize low-income beneficiaries who are at higher risk or older beneficiaries who have transportation barriers for healthcare access Nelson et al. (2016); Syed et al. (2013) via the following preference prompt: Prioritize low-income beneficiaries and older beneficiaries.
@@
Unfortunately, handcrafting reward functions is often a challenging and time-consuming task for humans because of the complex relationship between reward functions and policy outcomes Hazra et al. (2024); Cao et al. (2024); Ma et al. (2023). However, the multiagent nature of the RMAB problem adds a new twist to the problem of reward design in RL: It becomes fundamentally multi-objective. Consider the above example prompt asking for the prioritization of two subpopulations. As these subpopulations may contain different agents, selecting a reward function will most likely involve trading off the interests of the low-income vs. older beneficiaries, making this a multi-objective problem. If this multi-objective nature is ignored, a selected reward function might heavily favor one of the two groups (e.g., leading to the allocation of many resources to low-income beneficiaries, and no resources to older ones). Another tradeoff that emerges because of the multiagent nature of the problem is between aligning with the preference prompt and impacting agents not referenced in the prompt: Changing the reward functions to align with a prompt can lead to drastic changes in the number of resources distributed to subpopulations not referenced in the user prompt (e.g., in the example prompt, the selected reward might lead to taking away all resources from well-educated beneficiaries). If not adequately balanced, this can cause strong side effects unintended by the user.
@@
We present a Social Choice Language Model (SCLM) that designs reward functions (as Python code) aligned with complex, multi-objective human language preferences; see Figure 3 for an overview of SCLM. Our pipeline separates the generation of candidate reward functions in the generator from the selection of one function in the adjudicator. For the generator, we use LLM-powered evolutionary search to generate a pool of reward functions Behari et al. (2024). In the transparent and customizable adjudicator, we take a new social choice perspective to address the multi-objective nature of our problem: We create a scorer component that evaluates the quality of generated reward functions according to the different objectives (e.g., different prioritization requests). Subsequently, a social welfare function aggregates these “alignment scores” to select the best reward function. The user can select the social welfare function and thereby has additional control over the preferred trade-off between objectives, e.g., maximizing the summed vs. minimum alignment of all objectives. We demonstrate in our experiments that our SCLM leads to the selection of reward functions significantly better aligned with complex, multi-objective prompts. Moreover, we also show how it can be used to effectively mitigate the risks of using rewards designed from human prompts: unintended effects for other agents and the ineffective allocation of resources. SCLM combines the generative power of LLMs to design reward functions with the capabilities of social choice to handle multi-objective decision-making scenarios.
2. Related Works
@@
LLMs have emerged as a powerful tool to enhance RL. Recent work has used LLMs to generate reward functions based on natural language descriptions Ma et al. (2023); Xie et al. (2024b); Yu et al. (2023). For instance, Goyal et al. (2019); Carta et al. (2022); Mirchandani et al. (2021); Hazra et al. (2024) shape rewards by training an RL agent to learn and complete intermediate tasks guided by language, yet focusing on very different (non-multiagent) environments.
The work of Behari et al. (2024) is the first to present a Decision-Language Model for generating reward functions for RMABs from human prompts. The model performs a form of evolutionary search to find reward functions aligned with the given prompt in two interleaving phases: generation and reflection. In the generation phase, an LLM generates a set of reward functions. Based on reward function’s performances, in the reflection phase Ma et al. (2023); Shinn et al. (2023), the LLM selects the function best aligned with the prompt. This function is then included in the prompt for the next generation phase or returned. In contrast to our work, DLM mixes generation with selection and does not explicitly account for the multi-objective nature of the reward design problem. Furthermore, in contrast to our work, they focus on small RMAB instances ($\sim 20$ arms). Throughout the paper, we will use a slightly modified variant of DLM adjusted to our setting (see Section B.2) as a baseline.
LLMs can also be integrated into model-based and explainable RL, serving as world model simulators Lin et al. (2023) or policy interpreters Das et al. (2023). They can also act as information processors, effectively managing environment data and task instructions Basavatia et al. (2024); Song et al. (2023) and as direct decision makers Li et al. (2022) to improve the sample efficiency of offline RL. While this area of research is rapidly evolving, there remains a significant gap in addressing multi-objective, multi-agent scenarios within LLM-enhanced RL, which we begin to address in this paper.
@@
Research on MORL focuses on learning policies that maximize (and balance between) multiple objective functions, typically via scalarizing the objectives into a single reward function Moffaert et al. (2013) or approximating the Pareto front Roijers et al. (2013); Van Moffaert and Nowé (2014). In the context of multiagent systems, MORL has been used as a method to ensure the fair treatment of the individual agents Jiang and Lu (2019); Zimmer et al. (2021); Zimeng et al. (2023). Closest to ours from this line of work is the paper by Zimeng et al. (2023), who use ideas from the resource allocation literature to combine multiple objectives into a singular non-linear objective function. However, in contrast to our paper, their focus lies on policy learning for such non-linear objective functions. They neither consider reward design, human preference prompts, LLMs, nor RMABs.
Inverse Reinforcement Learning (IRL)
Another area related to this work is inverse reinforcement learning (IRL) Arora and Doshi (2021); Jain et al. (2024), which also focuses on obtaining the reward function. However, IRL aims to recover the underlying reward function that drives an expert’s demonstrated behavior. In other words, IRL seeks to understand why an agent acts in a certain way by observing its actions, rather than directly specifying its goals. In contrast, our approach does not depend on expert demonstrations, as such data is unavailable in our setting. Instead, we explicitly optimize the reward function to guide the agent toward the desired behavior.
@@
3. Preliminaries
An instance of Restless Multi-Armed Bandits (RMAB) is defined by a set of $N$ arms, a time horizon $T$, and a budget $K$. We also refer to arms as agents. Each arm $i\in[N]$ is an independently evolving MDP with state space $\mathcal{S}_{i}$, actions $\mathcal{A}_{i}=\{0,1\}$, transition function $P_{i}:\mathcal{S}_{i}\times\mathcal{A}_{i}\times\mathcal{S}_{i}\to\mathbb{R}_{\geq 0}$, and reward function $R_{i}:\mathcal{S}_{i}\to\mathbb{R}$. We refer to $1$ as the active action corresponding to pulling the arm (i.e., allocating a resource) and 00 as the passive action corresponding to not pulling the arm. We focus on the popular case where each MDP consists of two states, i.e., $\mathcal{S}_{i}=\{0,1\}$ for all $i\in[N]$, yet our methodology applies to MDPs with arbitrary state spaces. We refer to 00 as the bad and $1$ as the good state. For each step in which an agent is in the good state, they derive a utility of $1$, while they derive a utility of 00 in the bad state. Accordingly, agents’ default reward function $R^{*}$ is $R^{*}(s)=s$. We assume that there is a set of categorical features. Each arm is associated with a value of each feature. A global reward function is a reward function defined over features, which induces a reward function for each arm by plugging in its feature values (see Section 4).
In each step within the time horizon $T$, the planner observes the state of all arms and decides to pull a subset of at most $K$ arm. As solving the RMAB problem optimally is computationally intractable Papadimitriou and Tsitsiklis (1994), we make use of the very popular state-dependent Whittle index Whittle (1988); Niño-Mora (2023), which tries to quantify for each state of each arm the reward gain achieved from applying the active action to the arm in this state. In the Whittle index policy $\Pi$, in each step, we compute the Whittle index for each arm (based on its current state) and pull the arms with the $K$ highest Whittle indices. We will use it as the solution strategy in the following.
For a global reward function $R$, we write $\Pi(R)$ to denote the Whittle index policy for $R$, i.e., the Whittle index policy for the instance where each agent uses the function $R$ after plugging in their feature values as their reward. We refer to $\Pi(R^{*})$ as the default policy. To assess the quality of a global reward function $R$, we often consider the utility feature distribution for some feature $X$. This distribution shows for each value of the feature, the expected utility generated by arms with this feature value under the policy $\Pi(R)$ (see Figure 2(a) for an example).
4. Problem Statement & Challenges
@@
Figure 1. Tradeoffs between prioritization clauses.
@@
Another shortcoming of DLM are unintended utility shifts.
(a) Income feature
(b) Education feature
Figure 2. Utility feature distributions for the default reward function (orange) and reward function returned for prompt “Prioritize agents with low income” (blue). The $x$-axis depicts the values of the feature and the $y$-axis the total utility generated by all agents with this feature value.
Moving from the default reward function to a reward function aligned with a given (prioritization) prompt causes shifts within the distribution of resources and utility. Due to correlations between features, this change might lead to unintended utility shifts for features not specified in the prompt. Figure 2 shows an example of this from our experiments. We present the utility feature distribution for the two features income and education for two reward functions: The reward function selected by DLM for the prompt “Prioritize agents with low income” (orange) and the default reward (blue). While the utility generated by low-income agents increases when moving from the default to the customized reward function, the utility generated by highly educated agents decreases, a side-effect the end-user might be unaware of and that might conflict with their allocation goals. In our proposed approach, we are able to account for this issue by incorporating the prevention of unintended utility shifts as a tradeoff dimension.
Thus, through Social Choice Language Model: Our goal is to create a model that handles multiple tradeoffs posed by composite “multi-objective” prompts in a principled, transparent, and customizable fashion and outputs a single effective and fairly aligned global reward function.
5. Social Choice Language Model (SCLM)
Figure 3. Overview of SCLM. In step 1, the preference prompt is passed to the generator, which performs an evolutionary search to create a pool $\mathcal{R}$ of candidate reward functions. In step 2, these functions are passed to the adjudicator where a scorer model (e.g., the simulator or LLM scorer) computes the alignment scores. In step 3, a user-defined social welfare function selects a reward function based on the alignment scores.
@@ 5. Social Choice Language Model (SCLM)
5.1. Generator
Given a prompt, our generator creates a set of candidate reward functions (as Python code) via a variant of evolutionary search Behari et al. (2024): We proceed in multiple steps. First, inputting the problem description, feature descriptions and the preference prompt, we ask an LLM to generate code for a reward function. We repeat this query $n_{p}$ times to obtain a set $\mathcal{R}$ of $n_{p}$ candidate reward functions. Afterwards, for each function $R\in\mathcal{R}$ we compute the utility feature distributions of the policy $\Pi(R)$ induced by the reward function $R$ on the given RMAB instance (via repeatedly simulating the policy on the instance). Then, the prompt and the set of candidate reward functions together with the associated utility feature distributions are passed to an LLM, which is asked to select the reward function $R^{\prime}$ from $\mathcal{R}$ best aligned with the prompt Ma et al. (2023); Shinn et al. (2023). Now, we repeat the whole process, this time including the selected policy $R^{\prime}$ as a seed in the reward function generation prompts. Once we have executed the process $n_{r}$ times, we add all generated $n_{p}\cdot n_{r}$ candidate reward functions $R$ to the pool $\mathcal{R}$ (see Section B.2 for details).
5.2. Adjudicator
The adjudicator selects a reward function from a given pool of candidate reward functions returned by the generator. To handle the complex tradeoffs arising within composite prompts and the resulting multi-objective optimization problem, the adjudicator follows a social choice approach. Social choice is a discipline at the intersection of economics, philosophy, and mathematics and concerned with aggregating the potentially contradicting preferences of a set of voters into a fair compromise alternative from a given candidate set Arrow et al. (2010); Moulin (2004). It thus provides a theoretically grounded and well-studied methodology for balancing competing interests. In our problem, we can interpret the reward functions as the candidates and the preference clauses in the prompt as the voters with their preferences over the candidates reflecting the reward function’s alignment with the clause. This view gives rise to the following strategy: Given a prompt $P=\{p_{1},p_{2},\dots,p_{\ell}\}$, we evaluate each reward function $R\in\mathcal{R}$ from the perspective of each preference clause $p_{i}$ by computing an (alignment) score $s_{i}(R)$. $s_{i}(R)$ measures the alignment of $\Pi(R)$ with preference clause $p_{i}$, i.e., how much the voter representing $p_{i}$ “likes” the candidate $R$. We describe at the end of this section how these scores are computed.